Tags:
create new tag
view all tags

20101109 Shift

17:00 - 25:00 AA, LC, LD, LN

Plan

  • Access to check

  • OTR Camera
  • Remove card in front of OTR lens
  • Remove OTR Camera Cap
  • Check APD on

  • Reference wire scanners to alignment laser measurements previously
  • Reference wire scanners to BPMS
  • Load our laserwire 'compromise' optics
  • Compare to normal ATF2 optics for background and orbit
  • Change pointing of electron beam if needed to be pointing at our window
  • Minimise background
  • Investigate size at LW-IP with OTR - not possible as our OTR camera not working

Access

NB SLAC OTRs are not working just now either and they said not to play with them. OTR camera could not be found on network. Has done this previously. Tried for 30mins - nothing. Perhaps replace with GigE camera on Thursday.

Break

Teranuma-san tuned accelerator / laser for stability for a bit. Lost about 1:30hrs.

Referencing

Wire Scanner Dimension E-beam Scan Range Laser Intersection (mm)
MW1X X 38-40 23.3995
Y 68-70 52.4980
MW2X X 36-38 37.1490
Y 66-68 61.2980

Measurements to within 100 microns.

MW2X -1.15m - LWIP - 1.85m - MW1X

Detector - ~19m - LWIP

Window 27mm diameter, 13.5mm radius. Signal beam is around 1cm in diameter ("LD"). Say we want to be ±5mm at the window from our vector we have defined with the laser referencing.

Uncertainty from laser referencing measurements -> s=r*θ. With 200 microns from the two measurements separated by 3.00m -> θ=66 micro-radians. At a distance of say 20m this corresponds to 1.32mm.

Tolerance is ±5mm at 20m or 10mm in total at 20m -> 500 micro-radians. This corresponds to a maximum allowable difference in the wire scanner positions of 1.5mm! This is quite big!

Results

Wire Scanner Dimension E-beam Scan Range Laser Intersection (mm) Wire Scanner Intersection (mm) Difference σ (μm)
MW1X X 38-40 23.3995 21.861 1.5385 184.1±27.1
Y 68-70 52.4980 53.016 0.518 21.4±2
MW2X X 36-38 37.1490 34.486 2.663 54±2.8
Y 66-68 61.2980 61.585 0.287 24.1±2.9

This corresponds to an angle of 374μrad in X and 77μrad in Y. These correspond to a displacement of 7.5mm in X and 1.5mm in Y.

BPM Referencing

100 pulses were logged using lwLog.py. Data is in extlw/dat/raw/lwLog_20101109_203939.dat. This was taken without the wire from the wire scanner in the beam.

Load Laserwire Compromise Optics

Optics loaded (~21:00)

Background was ~2500 adc counts with ATF2 optics after Mark Woodley tuned down Plic monitor for us as it was much higher before that.

With our optics it started at ~12000. Checked wire scanners were in stay positions which helped.

Investigated various magnets against background.

Tuned down to ~6000 with quads.

First data were logged with normal optics (9th November 2010 20:37). Then compromise optics were loaded (1st June 09 15:33) one magnet at a time and data logged (including Cerenkov detector and charge in the extraction line, upstream and downstream) after each change in order to investigate various magnets against background.

In compromise optics, two quads upstream of the LWIP (QD18X and QF19X) are changed in order to focus the beam at the LWIP. The downstream quads are changed in order to recollimate. In some cases the magnet settings were tweaked and differ from the 1st June optics saved file. The new settings are shown in column "tweak" below. The format for the data files is /userhome/nevay/extlw/dat/raw/lwLog_20101109_HHMMSS.dat where HHMMSS is the time.

Results

Change Tweak Data file time stamp
QD18X and QF19X   204649
QD20X and QF21X   204906
QM16FF 9.216A (8.216A in file) 205227
QM15FF 22.220A (23.220A in file) 205638
QM14FF   210813
QM13FF 45.780A (36.780A in file) 211307
QM12FF   211835
QM11FF   212225

The remaining magnets in the final focus do not differ significantly from the saved file and are left unchanged.

From the BPMS we can determine that much of the beam loss is occuring at QF9A due to the small signal on the BPM there (see Lawrence's lab book for print out of screen shot showing BPM rfDebug.edl).

A lot of steering made a great deal of difference.

Changing ZH1X from 0 to 0.01A resulted in a drop of around 700 ADC counts.

ZH8 and 9 were used as dog leg to change the position of the beam which helped too.

Background of around 2500 - 3000 ADC counts. 100 pulses logged in lwLog_20101109_221725.dat file in extlw/dat/raw/

Background

Background was reduced to around 2300 ADC Counts which is what it was with normal ATF2 optics but now with our compromise optics. Unplugging from the ADC, it was ~1.5V which is still very high.

Wire Scanner Referencing

Wire scans were taken to determine where the beam was with our compromise optics and the direction of beam.

Wire Scanner Dimension Laser Intersection (mm) WS Normal Optics (mm) σ (μm) WS LW Compromise Optics (mm) σ (μm)
MW1X X 23.3995 21.861 184.1±27.1 20.731 168.3±11.7
Y 52.4980 53.016 1.4±2 53.158 16.4±1
MW2X X 37.1490 34.486 54±2.8 32.386 144.9±22.1
Y 61.2980 61.585 24.1±2.9 62.257 20.6±1.6

Before MW1X X scan - lwLog_20101109_225158.dat Before MW1X Y scan - lwLog_20101109_230438.dat Before MW2X X scan - lwLog_20101109_231136.dat

Data taken before the wire enters the beam and not during.

Scan of MW2X Y recorded in lwLog_20101109_233253.dat to look at profile of wire scanner.

These correspond to angles of

X 698μrad Y 99μrad

These correspond to a displacement of 13.9mm in X and 1.9mm in Y.

June Optics

Loaded quads upstream of us from 'set10jun01_1533.dat' and background dropped from ~2500 to ~1100. Pedestal was ~410. Checked voltage of PMT on scope and it was 20mV! Very good indeed.

This entire setup was saved as 'set10nov10_0004.dat'

Logged 100 pulses at this setup - saved in lwLog_20101110_001355.dat

Wire Scanner Reference with New Optics

Repeated the wire scan measurements to find out the pointing vector with these optics.

Wire Scanner Dimension Centre Width Data File
MW1X X 20.552 152.9±11.4 none
Y 53.109 12.7±0.7 lwLog_20101110_003643.dat
MW2X X 32.610 146±8.2 lwLog_20101110_012040.dat
Y 62.097 17.1±1.1 lwLog20101110_011145.dat

Edit | Attach | Watch | Print version | History: r3 < r2 < r1 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r3 - 15 Nov 2010 - LauraCorner

 
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback